OpenStudy (anonymous):

Determine if the following sequence converges or diverges:

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (-1)^n/n!\]

OpenStudy (anonymous):

ok, well if a series converges absolutely, that is \[\sum_{n=1}^{\infty} |(-1)^n/n!|\] converges then the original series converges as well.

OpenStudy (anonymous):

so does \[\sum_{n=1}^{\infty}| (-1)^n/n!|=\sum_{n=1}^{\infty} \frac{1}{n!}\] converge?

OpenStudy (anonymous):

1/e and e.

OpenStudy (aravindg):

ya